OpenStudy (anonymous):

The measure of an angle is 12° less than twice the measure of its supplement. What is the measure of the angle? 116° 28° 64° 34°

OpenStudy (anonymous):

x+y= 180 (Supplementary angles) Therefore, x= 180 - y Now, using the given information : x = 2( 180 - x) - 12 x= 360 - 2x -12 x + 2x = 360 - 12 3x = 348 Therefore x = 116. Hence the answer is a] 116

OpenStudy (anonymous):

is it 116° or 28° or 64° or 34°

OpenStudy (anonymous):

It is 116.

OpenStudy (anonymous):

Which statement can be combined with its converse to form a true biconditional? If a ray is the perpendicular bisector of a segment, then the ray divides the segment into two congruent segments. If an angle is a straight angle, then its sides are opposite rays. If the measure of an angle is 30, then it is an acute angle. If two lines intersect, then the two lines are not skew. wat one is this

OpenStudy (anonymous):

I think it's b] If an angle is a straight angle, then its sides are opposite rays. Because all the others' converses don't seem to be true. I'm a little doubtful about the last one though.

OpenStudy (anonymous):

Use the Distributive Property to complete the following statement: \[4(x-1)=4x-\] 1. –4. 4. –1.

OpenStudy (anonymous):

Which biconditional is NOT a good definition? A whole number is even if and only if it is divisible by 2. Two line segments are congruent if and only if they have same length. Three points are collinear if and only if they lie on the same line. Two lines are parallel if and only if they do not intersect.

OpenStudy (anonymous):

Use the Distributive Property to complete the following statement: 4(x−1)=4x− 1. –4. 4. –1.

OpenStudy (anonymous):

Read the following two statements. Then use the Law of Syllogism to draw a conclusion. If the TV is too loud, then it will give me a headache. If I have a headache, then I will have to rest. If the TV volume is too loud, then I will have to rest. If I rest, then the TV volume was too loud. If I have a headache, then the TV volume is too loud. If I rest, then I have a headache.